OpenStudy (anonymous):

Write the expression in the form bi, where b is a real number. The Square Root of -3

OpenStudy (anonymous):

\[i \sqrt{3}\]

OpenStudy (anonymous):

Is that the answer? If so, how do you figure it out?

OpenStudy (anonymous):

\[\sqrt{-3}=\sqrt{-1*3}=\sqrt{-1}*\sqrt{3}\] now \[\sqrt{-1}=i\] so you get \[i \sqrt{3}\]

OpenStudy (anonymous):

Hmm.. didn't know how to do that. Seems easy, Thanks :)

OpenStudy (anonymous):

welcome...☺